### v1

@@ 2.4 Decoding hand function
The following sections present the results of three offline analyses: 1) classification of all 16 classes of recorded tasks; 2) classification of a subset of 12 classes associated to tasks primarily requiring the muscles that were targeted by the multi-electrode arrays (see Section 2.1), and 3) classification of six (S1) and eight (S2) classes chosen as important for prosthetic control [15] (considered as a potential application of the proposed decoding system). In all analyses, results are reported when using three feature spaces: motor unit discharge timings as decoded from HD-iEMG, root mean square (r.m.s.) values for each channel of the HD-iEMG, and r.m.s. values for each channel of the HD-sEMG. The different classification pipelines are depicted in Figure 5. Additionally, in the first two analyses we report results from lower-density iEMG and sEMG signals (in Figure 6 b and d respectively). This analysis was performed to investigate whether high-density configurations were essential for achieving high prediction accuracy.
Figure 5: Analysis pipelines. a Classification using the MU discharge times as features. The EMG signals from each task (finger flexion and extension) were decomposed into individual motor unit activities. The motor unit action potential filters were then pooled into a filter bank, denoted as $\boldsymbol{\mathbb{F}}$. These filters were applied to the EMG signals of each task, generating a set of Innervation Pulse Trains (IPTs). Each IPT was clustered into source and noise components. If the clustering quality was robust, the source was accepted and segmented into sliding windows. Each time bin was assigned a class label based on the filter that produced the corresponding source. Majority voting was applied across the set of IPTs for each segment. The final classification label was determined by the ratio of correctly labelled segments to the total number of labelled segments. b Classification from global EMG features. The r.m.s. values of the EMG signals were extracted and segmented into sliding windows. These formed the input to a linear C-Support Vector Machine (SVM) classifier. The model was fit on the training data and then applied to the test data for classification. The resulting output provided the final classification labels.
@@ 2.4.1 Full classification (16 classes)
In the analysis of motor tasks spanning all 16 flexion and extension categories, the motor neuron discharge timings decoded from HD-iEMG provided a perfect classification accuracy for subject S1 of 100% and an accuracy of 96.1% for subject S2 on average across the tasks. The mis-classification for subject S2 was exclusively related to the little finger, and it was due to the fact that the electrode placement did not adequately target the EDM, the muscle primarily responsible for the activation of the little finger. The accuracy using the decomposition from the HD-iEMG recordings was higher than that achieved using the r.m.s. values of the HD-iEMG (90.1% and 94.3%) or the r.m.s. values of the HD-sEMG (96.6% for S1, 76.5% for S2). Figure 7 presents the confusion matrices per subject and classification type.
Figure 7: Normalised confusion matrices illustrating the percentage of predictions for each of the 16 flexion and extension gestures, comparing true labels (y-axis) with predicted labels (x-axis). MU stands for motor unit. The matrices show predictions for S1 on the left and S2 on the right, across different data types: a HD-iEMG motor units. b r.m.s. values of HD-iEMG. c r.m.s. values of HD-sEMG.
@@
When using low-density EMG configurations, the accuracy decreased substantially compared to the high-density systems, with the bipolar intramuscular achieving 47.8% for S1 and 69.0% for S2, and the sEMG achieving 92.5% for S1 and 64.3% for S2. These results indicate that the reduction in accuracy when shifting from a high-density to a lower-density configuration for pattern recognition is particularly large for intramuscular recordings and more modest for surface signals [16, 17]. The confusion matrices for the bipolar and low-density sEMG signals are shown in Appendix B, Figure 10.
@@ 2.4.2 Targeted muscle classification (12 classes)
Figure 8 shows the confusion matrices for the classes fully captured by the electrodes, as confirmed by ultrasound imaging (Fig. 2). The classifier using motor units decomposed from HD-iEMG signals reached 100% accuracy for both S1 and S2, which shows a perfect classification when targeting all muscle structures responsible for the tasks. The classifier using r.m.s. values from HD-iEMG achieved an accuracy of 99.3% for S1 and 99.7% for S2, while the classifier using r.m.s. values from HD-sEMG reached an accuracy of 95.3% for S1 and 71.4% for S2.
These results underscore the superior performance of intramuscular configurations over surface configurations for the classification based on neural features. When the electrodes accurately targeted the relevant muscles, the classification accuracy was 100% in the two subjects.
Figure 8: Normalised confusion matrices illustrating the percentage of predictions for each of the 12 flexion and extension gestures, comparing true labels (y-axis) with predicted labels (x-axis). MU stands for motor unit. The matrices show predictions for Subject S1 on the left and S2 on the right, across different data types: a HD-iEMG motor units. b r.m.s. values of HD-iEMG. c r.m.s. values of HD-sEMG.
@@
On average across the two subjects, the bipolar iEMG achieved an accuracy of 55.4% for S1 and 78.8% for S2, while the sEMG achieved an accuracy of 89.1% for S1 and 57.2% for S2. These results show a notable reduction in the accuracy compared to the high-density configuration systems, highlighting the importance of a high-density setup in achieving high classification accuracy. The confusion matrices for the 12-class classification from the lower-density configurations are shown in Appendix B, Figure 11.
@@ 2.4.3 Mapping specific functional classes (6 and 8 classes)
Figure 9 presents the classification results when focusing on the classes most relevant from a control perspective. Results are reported for six (S1) and eight (S2) classes (refer to Methods for details). The accuracy achieved from the motor unit activities decoded from HD-iEMG was 99.4% for S1 and 100% for S2, while the accuracy from the r.m.s. values of the HD-iEMG was 76.4% for S1 and 89.0% for S2. The accuracy from the r.m.s. values of the HD-sEMG was 87.0% for S1 and 74.4% for S2. The effectiveness of using high-density intramuscular EMG recordings combined with MU decomposition for neural prosthesis control is demonstrated also in this analysis by the very high accuracy rates (>99%). This near-perfect accuracy in classifying gestures typically used for prosthetic hand control indicates the potential of the proposed interface in external device control.
Figure 9: Normalised confusion matrix illustrating the percentage of predictions for each of the six (S1) and eight (S2) flexion and extension gestures, comparing true labels (y-axis) with predicted labels (x-axis). MU stands for motor unit. a motor unit HD-iEMG. b r.m.s. values of HD-iEMG c r.m.s. values of HD-sEMG.
@@ 3 Discussion
Decoding motor intent is critical for accurate control in neural interfacing systems. At the peripheral level, decoding performance depends heavily on the type and quality of recorded muscle signals. We developed an advanced intramuscular neural system capable of decoding the precise discharge timings of spinal motor neurons by deconvolving HD-iEMG recordings. These recordings capture detailed neural signals from motor neuron pools, offering rich insight into underlying motor control processes. The proposed system allowed us to observe motor neuron recruitment in single- and multi-digit tasks, revealing that a majority of motor units recruited during each task are task-specific, even when the same digits are used. These insights into motor neuron behavior suggest a strong potential for precise discrimination between hand tasks, which our system successfully validated with high accuracy.
Our recording and decoding system allowed us to explore the neural control during digit tasks. Notably, we observed that the motor units recruited during multi-digit tasks were almost entirely distinct from those recruited during the tasks of the corresponding single digits. This finding aligns with previous research by Oßwald et al. [18], which also suggested that motor unit recruitment during multi-digit tasks is well separated from that of single-digit tasks; however, this previous study was limited by the unconstrained and dynamic nature of the tasks investigated. Our results strongly indicate that hand control in multi-digit tasks is characterized by maximal separation of degrees of freedom, likely due to the demands of dexterous control. This was evidenced by the count of motor units active per task (Fig. 4). Additionally, we provided representative examples of the distinct action potential waveforms associated with different tasks, further supporting the highly specific recruitment of motor units during single- vs multi-digit tasks.
@@
Building on this understanding of the neural mechanisms underlying fine motor control, we hypothesized that a system leveraging the decoded activities of motor neurons could achieve near-perfect accuracy in gesture classification. This is because the detection of task-specific motor units is a perfect indicator of a task. The accuracy of the proposed mapping system was validated through a series of experiments. A novel signal classification method was applied to the decoded series of discharge timings, achieving near-perfect discrimination for six, eight and 12 classes (average >99%), and up to 16 classes (average >96%). While traditional HD-iEMG and HD-sEMG features demonstrated solid performance (over 71%), the motor neurons identified through HD-iEMG consistently yielded superior accuracy, exceeding global features by 0.3% to 28.6%, depending on the conditions. The substantial improvement in classification accuracy is highly relevant for applications in neuroprosthetics, as this increase in accuracy can significantly enhance the effectiveness and responsiveness of assistive devices, leading to better user experiences and outcomes. Moreover, in all tasks, the classification accuracy from HD-iEMG motor units for the combined finger movements approached 100% (Figs. 7 a, 8 a), further validating the task-specific recruitment of motor units during multi-digit tasks.
@@
Decoding motor intent can be approached through global EMG features or motor unit discharge timings, each with distinct advantages and challenges. Global EMG features represent the spatiotemporal convolution of many motor neuron discharges but are variable across individuals, particularly when recorded at the skin surface. This variability, alongside issues like amplitude cancellation and cross-talk, complicates precise control in complex tasks [19, 20]. Due to these challenges, global HD-sEMG achieved classification accuracies ranging from 71% to 97%, highlighting variability in performance across tasks.
@@
Although relatively high classification accuracies were achieved in some tasks, the reliance on features that are not directly linked to the underlying physiological mechanisms diminishes interpretability and presents challenges for clinical applications and motor control research. Understanding the specific contributions of muscle groups and neural activity is essential for validating model predictions and enhancing practical utility. In contrast to global EMG features, motor neuron discharge timings provide a direct and physiologically relevant representation of neural activity, closely reflecting the spinal-level coding mechanisms of movement. By decomposing the EMG signals into their constituent motor unit activities, we demonstrated that motor neuron spike trains can serve as critical sources of information for identifying complex tasks. This approach addresses the variability present in global EMG by separating neural information from muscle fiber action potentials, ensuring that the shapes of action potentials do not influence classification outcomes. Additionally, this separation mitigates the effects of EMG amplitude cancellation and enhances spatial and temporal sparseness, thereby reducing issues related to cross-talk [21].
@@
A key novelty of this work is the recording of high-density data intramuscularly, enabling detailed and accurate decoding of neural activity. The importance of intramuscular measurements is underscored by the large number of motor units that could be decomposed per task. High-density surface EMG has also been used for motor unit decoding [8] but it presents evident limitations with respect to implanted systems. One of these limitation is that it typically allows for the decoding of fewer motor units. This limitation is compounded by the challenge of accurately determining which muscles are activated, primarily due to a lack of spatial selectivity in the signals captured. This lack of resolution can lead to overlapping activity from adjacent muscles, making it difficult to pinpoint specific muscle contributions during movement [19, 20].
@@
Precise targeting of the appropriate muscles and decoding of direct neural information resulted in nearly perfect decoding accuracy from the motor unit activity derived from HD-iEMG. Ensuring accurate muscle targeting remains critical for effectively extracting useful information from these recordings. As we reported in Section 2.4.2, when there was high confidence in the targeted muscles, the accuracy of classification from decoded motor unit activities was 100%.
@@
Lower-density configurations were extracted from the high-density iEMG and sEMG data to evaluate whether multi-channel configurations were necessary for achieving high classification accuracies. The results showed that classification accuracy from global EMG features dropped with fewer channels, more significantly so for the iEMG. Regarding the 16 class classification, the bipolar iEMG configuration dropped by 40% for S1 and by 25% for S2 for 16 classes, while sEMG dropped by 4% for S1 and just over 10% for S2. A similar trend was observed for the 12 classes. This highlights the important of high-density configurations to accurately classify gestures from the EMG. It is important to note that classification based on motor neurons would not be feasible with lower-density configurations, as high-density setups are essential for the decomposition methods utilised to extract motor neuron activity.
@@
Various approaches have been explored in previous research for classifying hand gestures from global EMG features. For eight classes, time domain and deep learning methods have shown variable level of success, with accuracies ranging from 81.54% [22] to 95% and about 91% [23]. A convolutional auto-encoder combined with a CNN achieved 99% accuracy for 10 finger gestures [24], but this approach was tested on a smaller set of 10 classes compared to the 16 and 12 classes used in our study.
@@
There have also been previous studies that presented results on classification of motor unit activities, extracted from surface EMG, associated to hand tasks. For example, neural drives to muscle regions were estimated by grouping motor neurons based on the spatial location of their corresponding muscle units, achieving over 97% accuracy across 11 classes in patients with targeted nerve reinnervation [8], while similar methods using motor unit action potential waveform activations reached around 95% accuracy for 11 classes [25]. A related technique employing motor unit filters achieved 94.6% accuracy for 12 gestures [26]. Additionally, neuromorphic approaches have shown promise, with approximately 95% accuracy for 10 classes [27].
@@
Overall, our method outperforms all previous approaches, achieving accuracies exceeding 99% for six, eight, and 12-class classification tasks and greater than 96% for 16 classes. Notably, the precision obtained from motor unit activity significantly surpasses prior results, highlighting the robustness and reliability of our system for decoding neural information. The use of motor units not only enhances performance but also provides a theoretical framework for achieving 100% accuracy. As long as motor unit discharges are accurately detected, and each task involves at least one unique motor unit, it becomes theoretically possible to reach perfect classification. This implies that, with precise muscle targeting and continued advancements in decoding algorithms, flawless performance is achievable — a potential we have already demonstrated in some cases in this study (e.g., 100% accuracy for both S1 and S2 in the 12-class classification).
@@
Under the experimental conditions of our study, the proposed approach based on decoding motor neurons demonstrated superior accuracy compared to traditional pattern recognition methods. However, the small sample size, driven by the need for intramuscular insertions, limits the generalization of these results. Further validation with a larger population is needed.
@@
Transitioning to online implementation for real-world applications presents significant challenges. Real-time HD-iEMG decomposition is computationally intensive, requiring high sampling frequencies to capture neural signals accurately [28]. While advances in microprocessor speeds have made real-time neural data processing more attainable, the challenge remains significant due to the intensive computational demands, especially since even offline data recordings are often hindered by the high sampling rates required (this problem is clearly exacerbated when increasing the number of channels, as in this study).
@@
A promising future direction is the integration of spiking neural networks directly onto low power mixed-signal neuromorphic chips [29, 30]. As decomposition methods continue to advance and become more capable of adapting to signal nonstationarities [31], implementing real-time decomposition methods with neuromorphic electronic circuits will significantly accelerate the development of practical neural interfaces. The combination of adaptive signal processing and the computational efficiency of neuromorphic architectures represents a crucial step toward deploying neural prosthetics and other neurotechnologies that rely on real-time neural decoding [32]. This integration would not only enhance the system’s ability to handle dynamic, real-world scenarios but also reduce the need for constant retraining.
@@
Despite the clear advantages of intramuscular systems over non-invasive ones, practical challenges hinder their widespread adoption. Precise muscle targeting is essential, requiring careful planning and execution by trained personnel. Mis-targeting can lead to mis-classifications, while increasing the number of electrodes could improve targeting accuracy. Our study focused on well-controlled isometric contractions, where signal decomposition is highly reliable. The system remains untested in dynamic, naturalistic conditions where movement alters action potential shapes.
@@
In conclusion, we have described new findings on the motor control of multi-digit tasks. Additionally, we have proven the possibility of decoding the behaviour of the pools of motor neurons that physiologically innervate some of the muscles responsible for finger movements. This decoding was demonstrated in two healthy individuals with three implanted electrodes each and was shown to provide information that can be used for highly intuitive and accurate control commands. The accuracy of control derived from motor units extracted from HD-iEMG was near-perfect for both subjects and was successfully tested across a range of conditions, encompassing six to 16 distinct fine finger gestures.

### v2

@@ 2.3 Decoding hand function
The following sections present the results of three offline analyses: 1) classification of all 16 classes of recorded tasks; 2) classification of a subset of 12 classes associated to tasks primarily requiring the muscles that were targeted by the multi-electrode arrays (see Section 2.1), and 3) classification of six (participant one) and eight (participant two) classes chosen as important for prosthetic control (15) (considered as a potential application of the proposed decoding system). In all analyses, results are reported when using three feature spaces: motoneuron discharge timings as decoded from HD-iEMG (Fig. 4A, B), root mean square (RMS) values for each channel of the HD-iEMG, and RMS values for each channel of the HD-sEMG (Fig. 4C). Additionally, in the first two analyses, we report results from lower-density iEMG and sEMG signals (configurations visualised in Fig. 5B and D, respectively). This analysis was performed to investigate whether high-density configurations were essential for achieving high prediction accuracy.
Figure 4: Classification pipelines. A The HD-iEMG signals recorded by each electrode and task in the first recorded repetition were independently decomposed into their constituent neural activities. The separation filters optimised in the decomposition to extract the discharge times were pooled into a dictionary of filters. B Classification using the motoneuron discharge times as features. The separation filters were applied to the HD-iEMG signals of each electrode and task, in the second recorded repetition, generating a set of sequences of electrical pulses, also known as Innervation Pulse Trains (IPTs). Each IPT was clustered into source and noise components. If the clustering quality was robust, the source was accepted and segmented into sliding windows. Each time bin was assigned a class label based on the separation filter that produced the corresponding source. Majority voting was applied across the set of IPTs for each segment. The final classification label was determined by the ratio of correctly labelled segments to the total number of labelled segments. C Classification from global EMG features. The RMS values of the HD-EMG signals were extracted and segmented into sliding windows. These formed the input to a linear C-Support Vector Machine (SVM) classifier. The model was fit on the training data (repetition one) and then applied to the test data (repetition two) for classification. The resulting output provided the final classification labels.
@@ 2.3.1 Full classification (16 classes)
In the analysis of motor tasks spanning all 16 flexion and extension categories, the motoneuron discharge timings decoded from HD-iEMG provided a perfect classification accuracy for participant one of 100% and an accuracy of 96.1% for participant two on average across the tasks. The mis-classification for participant two was exclusively related to the little finger, and it was due to the fact that the electrode placement did not adequately target the EDM, the muscle primarily responsible for the activation of the little finger. The accuracy using the decomposition from the HD-iEMG recordings was higher than that achieved using the RMS values of the HD-iEMG (90.1% and 94.3%) or the RMS values of the HD-sEMG (96.6% for participant one, 76.5% for participant two). Figure 6 presents the confusion matrices per participant and classification type.
Figure 6: Normalised confusion matrices illustrating the percentage of predictions for each of the 16 flexion and extension tasks, comparing true labels (y-axis) with predicted labels (x-axis). The matrices show predictions for participant one on the first row and participant two on the second row, across different data types: HD-iEMG motoneurons (first column), RMS values of HD-iEMG (second column), RMS values of HD-sEMG (third column).
@@
When using low-density EMG configurations, the accuracy decreased substantially compared to the high-density systems, with the bipolar intramuscular achieving 47.8% for participant one and 69.0% for participant two, and the sEMG achieving 92.5% for participant one and 64.3% for participant two. These results indicate that the reduction in accuracy when shifting from a high-density to a lower-density configuration for pattern recognition is particularly large for intramuscular recordings and more modest for surface signals (16, 17). The confusion matrices for the bipolar and low-density sEMG signals are shown in Appendix B, Figure 9.
@@ 2.3.2 Targeted muscle classification (12 classes)
Figure 7 shows the confusion matrices for the classes fully captured by the electrodes, as confirmed by ultrasound imaging (Fig. 2). The classifier using motoneurons decomposed from HD-iEMG signals reached 100% accuracy for both participant one and participant two, which shows a perfect classification when targeting all muscle structures responsible for the tasks. The classifier using RMS values from HD-iEMG achieved an accuracy of 99.3% for participant one and 99.7% for participant two, while the classifier using RMS values from HD-sEMG reached an accuracy of 95.3% for participant one and 71.4% for participant two.
These results underscore the superior performance of intramuscular configurations over surface configurations for the classification based on neural features. When the electrodes accurately targeted the relevant muscles, the classification accuracy was 100% in the two participants.
Figure 7: Normalised confusion matrices illustrating the percentage of predictions for each of the 12 flexion and extension tasks, comparing true labels (y-axis) with predicted labels (x-axis). The matrices show predictions for participant one on the first row and participant two on the second row, across different data types: HD-iEMG motoneurons (first column), RMS values of HD-iEMG (second column), RMS values of HD-sEMG (third column).
@@
On average across the two participants, the bipolar iEMG achieved an accuracy of 55.4% for participant one and 78.8% for participant two, while the sEMG achieved an accuracy of 89.1% for participant one and 57.2% for participant two. These results show a notable reduction in accuracy compared to high-density configuration systems, highlighting the importance of a high-density setup in achieving high classification accuracy. The confusion matrices for the 12-class classification from the lower-density configurations are shown in the Appendix B, Figure 10.
@@ 2.3.3 Mapping specific functional classes (6 and 8 classes)
Figure 8 presents the classification results when focusing on the classes most relevant from a control perspective. Results are reported for six (participant one) and eight (participant two) classes (refer to Methods for details). The accuracy achieved from the motoneuron activities decoded from HD-iEMG was 99.4% for participant one and 100% for participant two, while the accuracy from the RMS values of the HD-iEMG was 76.4% for participant one and 89.0% for participant two. The accuracy from the RMS values of the HD-sEMG was 87.0% for participant one and 74.4% for participant two. The effectiveness of using high-density intramuscular EMG recordings combined with motoneuron decomposition for the neural control of prosthesis is also demonstrated in this analysis by the very high accuracy rates (>99%). This near-perfect accuracy in classifying tasks typically used for prosthetic hand control indicates the potential of the proposed interface in external device control.
Figure 8: Normalised confusion matrix illustrating the percentage of predictions for each of the six (participant one, first row) and eight (participant two, second row) flexion and extension tasks, comparing true labels (y-axis) with predicted labels (x-axis) across different tasks: motoneurons from the HD-iEMG (first column), RMS values of HD-iEMG (second column), and RMS values of HD-sEMG (third column).
@@ 3 Discussion
Decoding motor intent is critical for accurate control in neural interfacing systems. At the peripheral level, decoding performance depends heavily on the type and quality of recorded muscle signals. We developed an advanced processing system capable of decoding the precise discharge timings of spinal motoneurons by deconvolving HD-iEMG recordings. These recordings capture detailed neural signals from motoneuron pools, offering rich insight into the underlying motor control processes. The proposed system enabled precise discrimination between hand tasks, which was successfully validated with high accuracy.
@@
Given the high quality of the signals recorded from the intramuscular microelectrode arrays, we hypothesised that a system leveraging the decoded activities of motoneurons could achieve near-perfect accuracy in task classification. The accuracy of the proposed mapping system was validated through a series of experiments. A novel signal classification method was applied to the decoded series of discharge timings, achieving near-perfect discrimination for six, eight, and 12 classes (average >99%), and very high accuracy for up to 16 classes (average >96%). While traditional HD-iEMG and HD-sEMG features demonstrated solid performance (over 71%), the motoneurons identified through HD-iEMG consistently yielded superior accuracy, exceeding global features by 0.3% to 28.6%, depending on the conditions. The substantial improvement in classification accuracy is highly relevant for applications in neuroprosthetics, as this increase in accuracy can significantly enhance the effectiveness and responsiveness of assistive devices, leading to better user experiences and outcomes. Moreover, in all tasks, the classification accuracy from HD-iEMG motoneurons for the combined finger contractions approached 100% (Figs. 6, 7).
@@
Decoding motor intent can be approached through global EMG features or motoneuron discharge timings, each with distinct advantages and challenges. Global EMG features represent the spatiotemporal convolution of many motoneuron discharges but are variable across individuals, particularly when recorded at the skin surface. This variability, alongside issues like amplitude cancellation and cross-talk, complicates precise control in complex tasks (18, 19). Due to these challenges, global HD-sEMG achieved classification accuracies ranging from 71% to 97%, highlighting variability in performance across tasks. Although relatively high classification accuracies were achieved in some tasks, the reliance on features that are not directly linked to the underlying physiological mechanisms diminishes interpretability and presents challenges for clinical applications and motor control research. Understanding the specific contributions of muscle groups and neural activity is essential for validating model predictions and enhancing practical utility.
@@
In contrast to global EMG features, motoneuron discharge timings provide a direct and physiologically relevant representation of neural activity, closely reflecting the spinal-level coding mechanisms of movement. By decomposing the EMG signals into their constituent motoneuron activities, we demonstrated that motoneuron spike trains can serve as critical sources of information for identifying complex tasks. This approach addresses the variability present in global EMG by separating neural information from muscle fiber action potentials, ensuring that the shapes of action potentials do not influence classification outcomes. Additionally, this separation mitigates the effects of EMG amplitude cancellation and enhances spatial and temporal sparseness, thereby reducing issues related to cross-talk (20).
@@
A key novelty of this work is the recording of high-density data intramuscularly, enabling detailed and accurate decoding of neural activity. The importance of intramuscular measurements is underscored by the large number of motoneurons that could be decomposed per task. High-density surface EMG has also been used for motoneuron decoding (8) but it presents evident limitations with respect to implanted systems. One of these limitation is that it typically allows for the decoding of fewer motoneurons. This limitation is compounded by the challenge of accurately determining which muscles are activated, primarily due to a lack of spatial selectivity in the signals captured. This lack of resolution can lead to overlapping activity from adjacent muscles, making it difficult to pinpoint specific muscle contributions during movement (18, 19).
@@
Precise targeting of the appropriate muscles and decoding of direct neural information resulted in nearly perfect decoding accuracy from the motoneuron activity derived from HD-iEMG. Ensuring accurate muscle targeting remains critical for effectively extracting useful information from these recordings. As we reported in Section 2.3.2, when there was high confidence in the targeted muscles, the accuracy of classification from decoded motoneuron activities was 100%.
@@
Lower-density configurations were extracted from the HD-iEMG and HD-sEMG data to evaluate whether multi-channel configurations were necessary for achieving high classification accuracies. The results showed that classification accuracy from global EMG features dropped with fewer channels, more significantly so for the iEMG. Regarding the 16 class classification, the bipolar iEMG configuration dropped by 40% for participant one and by 25% for participant two for 16 classes, while sEMG dropped by 4% for participant one and just over 10% for participant two. A similar trend was observed for the 12 classes. This highlights the importance of high-density configurations to accurately classify tasks from the EMG. It has to be noted, however, that in this study, we compared each recording type to only a single low-density configuration, which may not fully generalise to all conditions. On the hand hand, we believe that this provides a representative baseline for evaluating the impact of electrode density on classification performance. Similar results have indeed been obtained by changing the low-density recording in pilot analyses (results not shown).
In a recent study  (21), we showed that both the number of identifiable motoneurons and the accuracy of their decomposition depend on the number and density of channels in intramuscular microelectrode arrays. Consequently, classification based on motoneuron activity would not be feasible with lower-density configurations, as the decomposition methods used to extract motoneuron signals require high-density setups. For this reason, we did not test decomposition-based features in the case of low-density recordings.
@@
Various approaches have been explored in previous research for classifying hand tasks from global EMG features. For eight classes, time domain and deep learning methods have shown variable level of success, with accuracies ranging from 81.54% (22) to 95% and about 91% (23). A convolutional auto-encoder combined with a CNN achieved 99% accuracy for 10 finger tasks (24), but this approach was tested on a smaller set of 10 classes compared to the 16 and 12 classes used in our study.
@@
There have also been previous studies that presented results on the classification of motoneuron activities, extracted from surface EMG, associated to hand tasks. For example, neural drives to muscle regions were estimated by grouping motoneurons based on the spatial location of their corresponding muscle units, achieving over 97% accuracy across 11 classes in patients with targeted nerve reinnervation (8), while similar methods using motoneuron action potential waveform activations reached around 95% accuracy for 11 classes (25). A related technique employing motoneuron filters achieved 94.6% accuracy for 12 tasks (26). Additionally, neuromorphic approaches have shown promise, with approximately 95% accuracy for 10 classes (27).
@@
Overall, our method outperformed all previous approaches, achieving accuracies exceeding 99% for six, eight, and 12-class classification tasks and greater than 96% for 16 classes. Notably, the precision obtained from motoneuron activity significantly surpasses prior results, highlighting the robustness and reliability of our system for decoding neural information. The use of motoneurons not only enhances performance but also provides a theoretical framework for achieving 100% accuracy. As long as motoneuron discharges are accurately detected, and each task involves at least one unique motoneuron, it becomes theoretically possible to reach perfect classification. This implies that, with precise muscle targeting and continued advancements in decoding algorithms, flawless performance is achievable — a potential we have already demonstrated in some cases in this study (e.g., 100% accuracy for both participant one and participant two in the 12-class classification). Reaching perfect, or almost perfect, accuracy in decoding is essential, particularly in clinical and real-world applications when even small errors can have significant consequences. For assistive devices, such as prosthetic limb control, near-perfect accuracy is often necessary to ensure reliable, intuitive, and safe operation. Errors in decoding can lead to unintended movements, frustration, or even device rejection by the user. While simpler systems may suffice for some applications, such as general activity recognition, tasks requiring fine motor control demand higher accuracy to maintain usability and effectiveness. The present results bring us substantially closer to this goal.
@@
Under the experimental conditions of our study, the proposed approach based on decoding motoneurons demonstrated superior accuracy compared to traditional pattern recognition methods. However, the small sample size, driven by the need for intramuscular insertions, limits the generalisation of these results. Further validation with a larger population is needed.
@@
Transitioning to online implementation for real-world applications presents significant challenges. Real-time HD-iEMG decomposition is computationally intensive, requiring high sampling frequencies to capture neural signals accurately (28). While advances in microprocessor speeds have made real-time neural data processing more attainable, the challenge remains significant due to the intensive computational demands, especially since even offline data recordings are often hindered by the high sampling rates required (this problem is clearly exacerbated when increasing the number of channels, as in this study).
@@
A promising future direction is the integration of spiking neural networks directly onto low power mixed-signal neuromorphic chips (29, 30). As decomposition methods continue to advance and become more capable of adapting to signal nonstationarities (31), implementing real-time decomposition methods with neuromorphic electronic circuits will significantly accelerate the development of practical neural interfaces. The combination of adaptive signal processing and the computational efficiency of neuromorphic architectures represents a crucial step toward deploying neural prosthetics and other neurotechnologies that rely on real-time neural decoding (32). This integration would not only enhance the system’s ability to handle dynamic, real-world scenarios but also reduce the need for constant retraining.
@@
Despite the clear advantages of intramuscular systems over non-invasive ones, practical challenges hinder their widespread adoption. Precise muscle targeting is essential, requiring careful planning and execution by trained personnel, as mis-targeting can lead to misclassifications. While increasing the number of electrodes may improve targeting accuracy, it also raises concerns about computational demands and real-time processing requirements. Additionally, intramuscular systems introduce substantial challenges beyond electrode placement. Their invasive nature increases the risk of infection, electrode migration, and potential tissue damage over time. Surgical implantation and the routing of leads present technical challenges, while transcutaneous connectors pose risks related to physical stability and long-term biocompatibility. It is also important to acknowledge that our study focused on well-controlled isometric contractions, where signal decomposition is highly reliable, but the system remains untested in dynamic, naturalistic conditions where movement alters action potential shapes.
@@
In conclusion, we have proven the possibility of decoding the behaviour of the pools of motoneurons that physiologically innervate some of the muscles responsible for finger contractions. This decoding was demonstrated in two healthy individuals with three implanted electrodes each and was shown to provide information that can be used for highly intuitive and accurate control commands. The accuracy of control derived from motoneurons extracted from HD-iEMG was near-perfect for both participants and was successfully tested across a range of conditions, encompassing six to 16 distinct fine finger tasks.